OpenStudy (anonymous):

Find intersection of the lines l1: r=(2,1,0) + t(1,-1,1) l2: r=(3,0,1) + s(2,3,-1)

OpenStudy (michele_laino):

Please try to re-write your lines with cartesian equations

OpenStudy (anonymous):

Set the vectors equal to each other:\[\begin{align*}\langle2,1,0\rangle+t\langle1,-1,1\rangle&=\langle3,0,1\rangle+s\langle2,3,-1\rangle\\ \langle2+t,1-t,t\rangle&=\langle3+2s,3s,1-s\rangle \end{align*}\] This gives the following system\[\begin{cases}2s-t=-1\\ 3s+t=1\\ s+t=1\end{cases}\]